# Estimation of uncertainties in the density driven flow in fractured porous media using MLMC

### v1

@@ 2.2 Model problem settings
For our numerical tests, we choose a simple but very illustrative two-dimensional problem with one fracture, proposed in [20]. It extends the setting of the Henry problem [25, 43] with a fracture. The aquifer is represented by a rectangular domain $\mathcal{D}=[0,2]\times[-1,0]$ $[\mathrm{m}^{2}]$ completely saturated with the liquid phase, see Fig. 1 (left). The fracture is located near the right “sea side” where the heavy salty water intrudes into the aquifer and cuts this boundary.
Figure 1: Left: Geometry, boundary conditions and the selected points for the modified Henry problem with a fracture. Right: The flow pattern of this problem; color: $c_{m}$ (red for $c_{m}=1$); black lines: streamlines of the flow field.
The model presented in Section 2.1 must be closed by specification of boundary conditions for $c_{m,f}$ and $p_{m,f}$, as well as initial conditions for $c_{m,f}$. In this work, we follow the settings from [20] but add the uncertainty as described in Section 2.3 below. In particular, for the initial conditions, we set
@@
where $\mathbf{e}_{x}=(1,0)^{\top}$, and $\hat{q}_{\mathrm{in}}$ is a prescribed function of time, see Sect. 2.3. Note that the fracture does not cut the left side of the domain, so no boundary conditions for $c_{f}$ and $p_{f}$ are required there. For the upper and the lower boundaries of $\mathcal{D}$, we impose no-flux boundary conditions for all the equations.
The flow and salt transport patterns appearing in this problem are discussed in [20]. They are presented in Fig. 1 (right): The color codes the mass fraction $c_{m}$ with red corresponding to $c_{m}=1$ and blue to $c_{m}=0$; the lines are the streamlines of the flow. In the left part of $\mathcal{D}$, the flow is induced by the pure water recharge through the left boundary. As in the Henry problem, the salt water intruding through the lower part of the right boundary turns up and is washed out through the upper part of the right side. In particular, the much more permeable fracture serves as a pathway for the strong wash-out that evacuates the salty water from the domain. For that, the part of the domain below the fracture is mostly separated from the upper part of the domain — as much as the capacity of the fracture allows. These phenomena depend on the width and the permeability of the fracture. In this work, we investigate how the fracture aperture influences the concentration.
In the case of time-independent boundary conditions, the mass fraction field converges to a steady state (cf. [20]). However, before it, there is an initial instationary phase taking considerable time when the maximum uncertainty in the solution w.r.t. the parameters is attained. This phase can be recognized even for the time-dependent recharge $\hat{q}_{\mathrm{in}}$ considered below.
@@
$\rho(c)=\rho_{0}+(\rho_{1}-\rho_{0})c,\qquad\qquad\mu=\text{const},$ (11)
where $\rho_{0}$ is the density of the pure water and $\rho_{1}$ is the density of the brine (both considered as constants). Besides that, we assume that porous medium is isotropic:
@@
${\mathbf{D}}_{m}=\phi_{m}D_{\mathrm{mol}}\mathbf{I},\qquad{\mathbf{D}}_{f}=\phi_{f}D_{\mathrm{mol}}$ (13)
where $D_{\mathrm{mol}}\in\mathbb{R}$ is the constant molecular diffusion coefficient of the salt in the liquid phase. For the interface between the bulk medium and the fracture, we assume
$K_{fn}=K_{m},\qquad D_{fn}=\phi_{m}D_{\mathrm{mol}},\qquad x\in\mathscr{S}.$ (14)
@@
We investigate the mass fraction $c_{m}$ at 6 locations where the largest variance is observed. This corresponds to the situation when the water quality of wells reaching these points is controlled. The list of selected points follows (see Fig. 1):
@@
Whereas the other points can be considered, the reason to select these 6 points (and 6 small subdomains) is that not all other points are “interesting”, i.e., not all points have significant variation in $c_{m}$. MLMC reduces the variance, but if the initial variance is small, there is no need to use MLMC.
Symbol Quantity Value Unit $D_{\mathrm{mol}}$ Deterministic diffusion coefficient in the medium $18.8571\cdot 10^{-6}$ [$\mathrm{m}^{2}\textrm{ }\mathrm{s}^{-1}$] $\mathbf{g}$ Gravity $9.8$ [m $\mathrm{m}\textrm{ }\mathrm{s}^{-2}$] $K_{m}$ Deterministic permeability of the medium $1.019368\cdot 10^{-9}$ [$\mathrm{m}^{2}$] $K_{f}$ Deterministic permeability of the fracture $1.019368\cdot 10^{-6}$ [$\mathrm{m}^{2}$] $\phi_{m}$ Deterministic porosity of the medium $0.35$ $\phi_{f}$ Deterministic porosity of the fracture $0.7$ $\mu$ Viscosity $10^{-3}$ [kg $\mathrm{kg}\textrm{ }\mathrm{m}^{-1}\textrm{ }\mathrm{s}^{-1}$] $\rho^{pW}$ Density of water $1\cdot 10^{3}$ [kg $\mathrm{kg}\textrm{ }\mathrm{m}^{-3}$] $\rho^{pB}$ Density of brine $1.025\cdot 10^{3}$ [kg $\mathrm{kg}\textrm{ }\mathrm{m}^{-3}$] $\epsilon$ Fracture width $\mathbb{E}\left[\epsilon\right]=5.05\cdot 10^{-3}$ [$\mathrm{m}$]
Table 1: Simulation parameters for the model problem from Section 2.2
@@ 2.3 Stochastic modeling of porosity, permeability and fracture width
Stochastic modeling is used to simulate the behavior of a fractured reservoir under uncertainty. The primary sources of uncertainty are the hydrogeological properties of the porous medium — porosity ($\phi_{m,f}$) and permeability ($K_{m,f}$) fields of the porous matrix and the fracture aperture ($\epsilon$) — as well as the freshwater recharge intensity $\hat{q}_{x}$. In this work we do not consider the uncertainty in $\phi_{f}$ and $K_{f}$. The QoIs are related to the mass fraction $c_{m}$, a function of $\phi_{m}$, $K_{m}$, and $\hat{q}_{x}$. We model the uncertain $\phi_{m}$ using a random field and assume $K_{m}$ to dependent on $\phi_{m}$:
@@
where $\kappa_{KC}$ is a constant scaling factor. The recharge inflow flux is kept constant across the left boundary but depends on the stochastic variable $\hat{q}_{x}$. We also assume that it is independent of $\phi$ and ${\mathbf{K}}$.
@@
The corresponding permeability is defined by (17).
Unknown porosity can be also modeled by a random field. One of the ways to compute this random field is to make an assumption about the covariance matrix or to estimate it from the available measurement data. Then an auxiliary eigenvalue problem is solved and the truncated Karhunen-Loéve expansion (KLE) is constructed [29]. Due to the high complexity of this approach, we implement a simpler one. We use $L_{2}$ orthogonal functions ($\sin()$ and $\cos()$), which mimic the $L_{2}$-orthogonal functions used in the KLE. These functions are then multiplied by the uniform random variables $\bm{\xi}$ as in (20). Typically one takes one of three options for $\bm{\xi}$: uniform, Gaussian or log-normal.
@@ 3.1 Numerical methods for the deterministic problem
The system (1)-(3) is numerically solved in the domain $\mathcal{D}\times[0,T]$. $\mathcal{D}$ is covered with an unstructured grid $\mathcal{D}_{h}$ of triangles and quadrilaterals. We denote the characteristic mesh size by $h$. We apply a special technique for resolving the jump of the solution at the fracture, cf. [36] for details. The vertex-centered finite-volume scheme is used for the discretization of the system (14) in space (cf. [36, 38]). The number of degrees of freedom associated with $\mathcal{D}_{h}$ is denoted by $n$. There are two degrees of freedom per grid vertex in $\mathcal{D}_{h}$: one for the mass fraction $c_{m,f}$ and another for the pressure $p$. At the fracture, several grid vertices share the same geometrical position, cf. [36, 38]. We use the regular refinement to create the grid hierarchy so that $h=\mathcal{O}(n^{-1/d})$. The implicit Euler method with a time step $\tau$ is used for the time discretization of (14). The number of the computed time steps is $r=T/\tau$. In this paper, we focus on the mass fraction $c_{m}$ of the salt in the bulk porous medium. We denote the part of the discrete solution of the model approximating $c_{m}$ by $c_{m,h,\tau}$.
@@
The implicit time-stepping scheme is unconditionally stable but requires solution of the extensive nonlinear discretized algebraic system with $n$ unknowns in every time step. The Newton’s method is used for this. Linear systems in the Newton iteration are solved using the BiCGStab method (cf. [2]) preconditioned with the geometric multigrid method (V-cycle, cf. [22]). In the multigrid cycle, the ILUβ-smothers [23] and Gaussian elimination are used as the coarse grid solver.
@@ 3.2 MLMC Algorithm
To reduce the total computing cost, we apply the MLMC method, which is a natural idea because the deterministic solver uses a multigrid method (see Section 3.1). The MLMC method gains efficiency by combining samples computed on different grids in the hierarchy. A more in-depth description of these techniques is found in [9, 10, 17, 18, 24, 47, 30].
Let $\bm{\xi}$ be a vector of random variables, and $g(\bm{\xi})$ the quantity of interest. In this work, we tried different QoI $g$: solution $c_{m}$ at a point $(t,\mathbf{x})$, integrals over a small subdomain $\Delta_{i}$ and over the whole computational domain $\mathcal{D}$, see Sect. 4. The MLMC method aims to approximate the expected value $\mathbb{E}\left[g\right]$ with an optimal computational cost. MLMC constructs a telescoping sum, defined over a sequence of spatial and temporal meshes, $\ell=0,\ldots,L$, as described next, to achieve this goal. The QoI $g$, numerically evaluated on level $\ell$, is denoted by $g_{h_{\ell},\tau_{\ell},\ell}$ or, for simplicity, by just $g_{\ell}$, where $h_{\ell}$ and $\tau_{\ell}$ are the discretization steps in space and time on level $\ell$.
Let $s_{0}$ denote the maximum computational cost of the evaluation of one realization of $g_{0}$. Similarly, $s_{\ell}$ denotes the computing cost of evaluating $g_{\ell}-g_{\ell-1}$. For simplicity, we assume that $s_{\ell}$ for $g_{\ell}-g_{\ell-1}$ is almost the same as $s_{\ell}$ for $g_{\ell}$. Since the number of required iterations is variable, the cost of computing a sample of $g_{\ell}-g_{\ell-1}$ may fluctuate for various realizations.
The MLMC method calculates $\mathbb{E}\left[g_{L}\right]\approx\mathbb{E}\left[g\right]$ using the following telescopic sum:
$\mathbb{E}\left[g_{L}\right]$ $=\mathbb{E}\left[g_{0}\right]+\sum_{\ell=1}^{L}\mathbb{E}\left[g_{\ell}-g_{\ell-1}\right]$ (22)
@@
In the above equation, level $\ell$ in the superscript $(\ell,i)$ indicates that independent samples are used at each correction level.
As $\ell$ increases, the variance of $g_{\ell}-g_{\ell-1}$ decreases. Thus, the total computational cost can be reduced by taking fewer samples on finer meshes.
In the following numerical experiments $n_{\ell}=4n_{\ell-1}=\ldots=4^{\ell}n_{0}=2^{d\ell}n_{0}$, $d=2$, and $r_{\ell}=2r_{\ell-1}=\ldots=2^{\ell}r_{0}$. In the case of uniform, equidistant mesh, we could also write similar formulas for step sizes: $h_{\ell}=h_{\ell-1}\cdot 2^{-1}=h_{\ell-2}\cdot 2^{-2}=\ldots=h_{0}\cdot 2^{-\ell}$ and $\tau_{\ell}=\tau_{0}\cdot 2^{-\ell}$.
The average cost $s_{\ell}$ of generating one sample of $g_{\ell}$ (the cost of one deterministic simulation for one random realization) is
$s_{\ell}=\mathcal{O}(n_{\ell}r_{\ell})=\mathcal{O}(4^{\ell}n_{0}\cdot 2^{\ell}r_{0})=\mathcal{O}(2^{2\ell}n_{0}\cdot 2^{\ell}r_{0})=\mathcal{O}(2^{\hat{d}\ell\gamma}n_{0}r_{0}),$ (24)
where $\hat{d}=d+1=3$ and $\gamma=1$.
@@
Denote by $Y:=\sum_{\ell=0}^{L}Y_{\ell}$ the multilevel estimator of $\mathbb{E}\left[g\right]$ based on $L+1$ levels and $m_{\ell}$ independent samples on level $\ell$, where $\ell=0,\ldots,L$.
@@
Using preliminary numerical tests (see Fig. 5) and following preprocessing steps as in Algorithm 1, we can estimate the convergence rates $\alpha$ for the mean (the so-called weak convergence) and $\beta$ for the variance (the so-called strong convergence), as well as the constants $C_{1}$ and $C_{2}$. In addition, $\alpha$ is strongly connected to the order of the discretization error (see Section 3.1), which equals $1$. Note that precise estimates of parameters $\alpha$ and $\beta$ are crucial to distribute the computational effort optimally.
@@
Input: $L=3$, $m_{\ell}=10$, $\ell=0,\ldots,L$
Compute $g_{\ell}(\omega_{i})$, $g_{\ell}(\omega_{i})-g_{\ell-1}(\omega_{i})$ for $i=1\ldots m_{\ell}$ and $\ell=1\ldots L$
Estimate $\mathbb{E}\left[g_{\ell}(\omega_{i})-g_{\ell-1}(\omega_{i})\right]$, $\mathbb{V}\left[g_{\ell}(\omega_{i})-g_{\ell-1}(\omega_{i})\right]$
Estimate convergence rates $\alpha$ and $\beta$, complexity rate $\gamma$, constants $C_{1},C_{2},C_{3}$
@@
Estimate $L$
Compute $m_{\ell}$, $\ell=0,\ldots,L$
Compute $\mathbb{E}\left[g_{\ell}(\omega_{i})-g_{\ell-1}(\omega_{i})\right]$, $i=1\ldots m_{\ell}$
Build the telescopic sum to estimate $\mathbb{E}\left[g\right]$

### v3

@@ 2.2 Model problem settings
For our numerical tests, we choose a simple but very illustrative two-dimensional problem with one fracture, proposed in [21]. It extends the setting of the Henry problem [27, 46] with a fracture. The aquifer is represented by a rectangular domain $\mathcal{D}=[0,2]\times[-1,0]$ $[\mathrm{m}^{2}]$ completely saturated with the liquid phase, see Fig. 1 (left). The fracture is located near the right “sea side” where the heavy salty water intrudes into the aquifer and cuts this boundary. The fracture coordinates are $(1,-0.7)$ and $(2,-0.5)$.
Figure 1: Left: Computational geometry $[0,2]\times[-1,0]$, direction of recharge, boundary conditions, location of fracture (marked by black oblique line) and six points, solution in which we use as additional QoIs. Right: The flow streamlines (thin black lines) and the salt mass fraction. The dark red colour indicates a high salt mass fraction $c_{m}=1.0$ or close to 1.0, and the dark blue corresponds to $c_{m}=0$ (no salt). The thick black curved line indicates the area where the recharge flow meets the density difference induced flow.
The system of the PDEs presented in Section 2.1 must be closed by specification of boundary conditions for $c_{m,f}$ and $p_{m,f}$, as well as initial conditions for $c_{m,f}$. In this work, we follow the settings from [21] but add the uncertainty as described in Section 2.3 below. In particular, for the initial conditions, we set
@@
where $\mathbf{e}_{x}=(1,0)^{\top}$, and $\hat{q}_{\mathrm{in}}$ is a prescribed function of time, see Sect. 2.3. For the upper and the lower boundaries of $\mathcal{D}$, we impose no-flux boundary conditions for Equations (1).
Remark 1.
For the model problem presented, we rescale the mass fractions $c_{m,f}$ so that $c_{m,f}=1$ corresponds to the most concentrated solution, i.e. seawater. Thus, seawater is considered as the abstract “salt”. This explains the condition (9) on the right boundary. Note that the density $\rho(c)$ is remapped accordingly, see (11). Therefore, the conservation equations for the salt (second equations in (1), (3) and (7)) are not affected.
The flow and salt transport patterns that occur in this problem are discussed in [21]. They are shown in Fig. 1 (right). The colour codes the mass fraction $c_{m}$, with dark red corresponding to $c_{m}=1$ and dark blue to $c_{m}=0$, and generally $c_{m}\in[0,1]$. The lines are the streamlines of the flow. In the left part of $\mathcal{D}$, the flow is induced by the pure water injection through the left boundary. As in the Henry problem, the saltwater intruding through the lower part of the right boundary turns up and is flushed out through the upper part of the right side. In particular, the much more permeable fracture acts as a pathway for the strong washout that removes the saline water from the domain. For this purpose, the part of the domain below the fracture is mostly separated from the upper part of the domain — as far as the capacity of the fracture allows. These phenomena depend on the width and permeability of the fracture.
In the case of time-independent boundary conditions, the mass fraction field converges to a steady state (cf. [21]). However, prior to this there will be an initial period of instability which may last for a considerable time. This phase can even be seen for the time-dependent recharge $\hat{q}_{\mathrm{in}}$ considered below.
@@
$\rho(c)=\rho_{0}+(\rho_{1}-\rho_{0})c,$ (11)
where $\rho_{0}$ is the density of the pure water and $\rho_{1}$ is the density of the brine (both considered as constants), and $\mu=\text{const}$. Besides that, we assume that porous medium is isotropic:
@@
${\mathbf{D}}_{m}=\phi_{m}D_{0}\mathbf{I},\qquad{\mathbf{D}}_{f}=\phi_{f}D_{0}$ (13)
where $D_{0}\in\mathbb{R}$ is the constant effective diffusion coefficient of the salt in the liquid phase. For the interface between the bulk medium and the fracture, we assume
$K_{fn}=K_{m},\qquad D_{fn}=\phi_{m}D_{0},\qquad x\in\mathscr{S}.$ (14)
@@
We examine the mass fraction $c_{m}$ at 6 locations where the largest variance is observed, i.e. we computed the variance in the whole domain and took locations where this variance is high. This corresponds to the situation when the water quality of wells reaching these points is controlled. The list of selected points follows (see Fig. 1):
@@
Whereas the other points can be considered, the reason to select these 6 points (and 6 small subdomains) is that not all other points are “interesting”, i.e., not all points have significant variation in $c_{m}$.
Symbol Quantity Value Unit $D_{\mathrm{mol}}$ Deterministic diffusion coefficient in the medium $18.8571\cdot 10^{-6}$ [$\mathrm{m}^{2}\textrm{ }\mathrm{s}^{-1}$] $\mathbf{g}$ Gravity $9.8$ [m $\mathrm{m}\textrm{ }\mathrm{s}^{-2}$] $K_{m}$ Deterministic permeability of the medium $1.019368\cdot 10^{-9}$ [$\mathrm{m}^{2}$] $K_{f}$ Deterministic permeability of the fracture $1.019368\cdot 10^{-6}$ [$\mathrm{m}^{2}$] $\phi_{m}$ Deterministic porosity of the medium $0.35$ $\phi_{f}$ Deterministic porosity of the fracture $0.7$ $\mu$ Viscosity $10^{-3}$ [kg $\mathrm{kg}\textrm{ }\mathrm{m}^{-1}\textrm{ }\mathrm{s}^{-1}$] $\rho_{0}$ Density of water $1\cdot 10^{3}$ [kg $\mathrm{kg}\textrm{ }\mathrm{m}^{-3}$] $\rho_{1}$ Density of brine $1.025\cdot 10^{3}$ [kg $\mathrm{kg}\textrm{ }\mathrm{m}^{-3}$] $\epsilon$ Fracture width $\mathbb{E}\left[\epsilon\right]=5.05\cdot 10^{-3}$ [$\mathrm{m}$]
Table 1: Simulation parameters for the deterministic model problem from Section 2.2
@@ 2.3 Stochastic modeling of porosity, permeability and fracture width
Stochastic modeling is used to simulate the behavior of a fractured reservoir under uncertainty. Among many possible sources of uncertainties, we consider only the following ones: hydrogeological properties of the porous medium — porosity ($\phi_{m}$) field of the surrounding porous matrix, the fracture aperture ($\epsilon$), as well as the freshwater recharge intensity $\hat{q}_{x}$. The permeability $K_{m}$ is considered to be dependent on the uncertain $\phi_{m}$. The QoIs we consider are the mean and variance of the mass fraction at the six observation points. We model the uncertain $\phi_{m}$ using a random field and assume $K_{m}$ to dependent on $\phi_{m}$:
@@
where $\kappa_{KC}$ is a constant scaling factor. In numerical simulations we used $\kappa_{KC}=1.5455\cdot 10^{-8}$ [$\mathrm{m}^{2}$].
The recharge inflow flux is kept constant across the left boundary but depends on the stochastic variable $\hat{q}_{\mathrm{in}}$. We also assume that it is independent on the porosities and the permeabilities in the domain.
@@
Other options for $\bm{\xi}$ would be Gaussian, log-normal, beta distributions. The corresponding permeability is defined by (17).
Remark 2.
The fracture model considered would not work for open (hollow) fractures. We assume that the fracture is filled with the same porous medium regardless of its opening. This is certainly a strong assumption. But under it, the porosity and (tangential and normal) permeabilities of the fracture should be independent of its aperture. Note that the aperture is considered to be uncertain, e.g. not accurately measured but not changing as in the case of poroelasticity. However, for the model presented, the effective permeability of the entire fracture increases with aperture: Thicker fractures induce faster flow, and we observe this in our calculations.
Unknown porosity can also be modeled by a random field. One of the ways to compute this random field is to make an assumption about the covariance matrix or to estimate it from the available measurement data. Then an auxiliary eigenvalue problem is solved and the truncated Karhunen-Loéve expansion (KLE) is constructed [31]. Due to the high complexity of this approach, we implement a simpler one. We use $L_{2}$ orthogonal functions ($\sin()$ and $\cos()$), which mimic the $L_{2}$-orthogonal functions used in the KLE. These functions are then multiplied by the uniform random variables $\bm{\xi}$ as in (20). Of course, the actual distribution may be different. Typically, one takes the uniform or Gaussian as an initial guess and then updates it using e.g. Bayesian update techniques. But this is already the inference/inverse problem. In addition, updating the distribution function requires measurement data, which we do not have.
@@ 3.1 Numerical methods for the deterministic problem
The system (13) is numerically solved in the domain $\mathcal{D}\times[0,T]$. $\mathcal{D}$ is covered with an unstructured grid $\mathcal{D}_{h}$ of triangles and quadrilaterals. We denote the characteristic mesh size by $h$. We apply a special technique for resolving the jump of the solution at the fracture, cf. [40] for details. The vertex-centered finite-volume scheme is used for the discretization of the system (14) in space (cf. [40, 42]). The number of degrees of freedom associated with $\mathcal{D}_{h}$ is denoted by $n$. There are two degrees of freedom per grid vertex in $\mathcal{D}_{h}$: one for the mass fraction $c_{m,f}$ and another for the pressure $p_{m,f}$. At the fracture, several grid vertices share the same geometrical position, cf. [40, 42]. We use the regular refinement to create the grid hierarchy so that $h=\mathcal{O}(n^{-1/d})$. The implicit Euler method with a time step $\tau$ is used for the time discretization of (14). The number of the computed time steps is $r=T/\tau$. In this paper, we focus on the mass fraction $c_{m}$ of the salt in the bulk porous medium. We denote the part of the discrete solution of the model approximating $c_{m}$ by $c_{m,h,\tau}$.
@@
The implicit time-stepping scheme is unconditionally stable but requires solution of the extensive nonlinear discretized algebraic system with $n$ unknowns in every time step. The Newton’s method is used for this. Linear systems in the Newton iteration are solved using the BiCGStab method (cf. [2]) preconditioned with the geometric multigrid method (V-cycle, cf. [23]). In the multigrid cycle, the ILUβ-smoothers [24] and Gaussian elimination are used as the coarse grid solver.
@@ 3.2 MLMC Algorithm
Let $\bm{\xi}$ be a vector of random variables, and $g(\bm{\xi})$ the quantity of interest (QoI). In this work, we tested different QoIs: solution $c_{m}$ at a point $(t,\mathbf{x})$, integrals $I_{i}$ over a small subdomain $\Delta_{i}$ and over the whole computational domain $\mathcal{D}$, see Sect. 4.
To estimate the propagation of uncertainties, we will calculate the mean and variance of the mass fraction. This statistical operation necessitates a large number of simulations, leading to high computational costs. The total cost can be approximated as the cost of a single simulation multiplied by the number of simulations.
To reduce the overall computational cost, we employ the well-known Multi-Level Monte Carlo (MLMC) method (see Algorithm 2). This approach is particularly suitable because the deterministic solver ug4 uses a multigrid method (see Section 3.1). The MLMC method enhances efficiency by combining samples computed on different grids within the hierarchy. A more detailed discussion of these techniques can be found in [9, 10, 18, 19, 26, 49, 32].
Below we provide an overview of the Multi-Level Monte Carlo (MLMC) method, drawing from various sources to ensure comprehensive coverage of implementation details. We also address aspects specific to our problem, including cost estimates at each level, convergence rates, and additional remarks. The MLMC method aims to approximate the expected value $\mathbb{E}\left[g\right]$ with an optimal computational cost. MLMC constructs a telescoping sum, defined over a sequence of spatial and temporal meshes, $\ell=0,\ldots,L$, as described next, to achieve this goal. The QoI $g$, numerically evaluated on level $\ell$, is denoted by $g_{h_{\ell},\tau_{\ell},\ell}$ or, for simplicity, by just $g_{\ell}$, where $h_{\ell}$ and $\tau_{\ell}$ are the discretization steps in space and time on level $\ell$.
Let $s_{0}$ denote the maximum computational cost of evaluating a realization of $g_{0}$. Similarly, let $s_{\ell}$ denote the computational cost of evaluating a realization of $g_{\ell}-g_{\ell-1}$. For simplicity, we assume that $s_{\ell}$ for $g_{\ell}-g_{\ell-1}$ is almost the same as $s_{\ell}$ for $g_{\ell}$. Since the number of iterations required is variable, the cost of computing a sample of $g_{\ell}-g_{\ell-1}$ may vary for different realizations.
Let $m_{\ell}$ be the number of random samples (numbers of realizations) on the $\ell$-th mesh. The MLMC method calculates $\mathbb{E}\left[g_{L}\right]\approx\mathbb{E}\left[g\right]$ using the following telescopic sum:
$\mathbb{E}\left[g_{L}\right]$ $=\mathbb{E}\left[g_{0}\right]+\sum_{\ell=1}^{L}\mathbb{E}\left[g_{\ell}-g_{\ell-1}\right]$
@@
In the equation above, the level $\ell$ in the superscript $(\ell,i)$ indicates that independent samples are used at each level. The samples used to sample $(g_{\ell}-g_{\ell-1})$ are different (independent) from the samples used to sample $(g_{\ell^{{}^{\prime}}}-g_{\ell^{{}^{\prime}}-1})$, for any $\ell^{{}^{\prime}}\neq\ell$. The upper index $(\ell,i)$ in the notation $(g_{\ell}^{(\ell,i)}-g_{\ell-1}^{(\ell,i)})$ means that the same sample is used for $g_{\ell}$ and $g_{\ell-1}$. There are $m_{0}$ independent samples for $g_{0}$, $m_{1}$ independent samples for $g_{1}-g_{0}$, etc. Thus, in total we generated $m_{0}+m_{1}+\cdots+m_{L}$ independent samples.
As $\ell$ increases, the variance of $g_{\ell}-g_{\ell-1}$ decreases. Thus, the overall computational cost can be reduced by taking fewer samples on finer meshes.
Remark 3.
To compute the second moment $\mathbb{E}\left[g^{2}\right]$, we proceed as follows, $\mathbb{E}\left[g_{L}^{2}\right]\approx\mathbb{E}\left[g^{2}\right]$ and then to get a telescopic sum we replace $g$ by $g^{2}$ in (22).
The information in this paragraph is problem specific. In the following numerical experiments $n_{\ell}=4n_{\ell-1}=\ldots=4^{\ell}n_{0}=2^{d\ell}n_{0}$, $d=2$, and $r_{\ell}=2r_{\ell-1}=\ldots=2^{\ell}r_{0}$. In the case of uniform, equidistant mesh, we could also write similar formulas for step sizes: $h_{\ell}=h_{\ell-1}\cdot 2^{-1}=h_{\ell-2}\cdot 2^{-2}=\ldots=h_{0}\cdot 2^{-\ell}$ and $\tau_{\ell}=\tau_{0}\cdot 2^{-\ell}$. The average cost $s_{\ell}$ of generating one sample of $g_{\ell}$ (the cost of one deterministic simulation for one random realization) is
$s_{\ell}=\mathcal{O}(n_{\ell}r_{\ell})=\mathcal{O}(4^{\ell}n_{0}\cdot 2^{\ell}r_{0})=\mathcal{O}(2^{2\ell}n_{0}\cdot 2^{\ell}r_{0})=\mathcal{O}(2^{\hat{d}\ell}n_{0}r_{0}),$ (23)
where $\hat{d}=d+1=3$.
@@
Let $Y:=\sum_{\ell=0}^{L}Y_{\ell}$ be the multilevel estimator of $\mathbb{E}\left[g\right]$ based on $L+1$ levels and $m_{\ell}$ be the number of independent samples at level $\ell$, where $\ell=0,\ldots,L$.
@@
The parameter $\gamma$ is the complexity rate of the numerical solver. Our geometric multigrid solver has linear complexity, i.e. $\gamma=1$.
@@
Using preliminary numerical tests (see Fig. 5) and following preprocessing steps as in Algorithm 1, we can estimate the convergence rates $\alpha$ for the mean (the so-called weak convergence) and $\beta$ for the variance (the so-called strong convergence), as well as the constants $c_{1}$ and $c_{2}$ from (35a)-(35b). In addition, the rate $\alpha$ is strongly related to the order of the discretisation error (see Section 3.1), which is equal to $1$. Note that precise estimates of the parameters $\alpha$ and $\beta$ are crucial for the optimal complexity of the MLMC method and for the optimal distribution of the computational effort over $L+1$ layers. The number of samples $m_{\ell}$ and the level $L$ will be suboptimal if the rates $\alpha$ and $\beta$ are not accurately estimated. To obtain these estimates, we run the Algorithm 1. Some advanced implementations of MLMC, e.g. CMLMC [10], can update the convergence rate on the fly. The calculated rates are used in the Theorem 2 and the Algorithm 2. This Algorithm 2 estimates the mean value of the QoI. Examples of calculated and used parameters $m_{\ell}$ and $L$ are listed in the rows of Table 4. The values in Table 4 are only valid for a fixed QoI. Both algorithms should be rerun, and all estimates made again for a new QoI.
@@
Input: set, for example, $L=3$, $m_{\ell}=10$, $\ell=0,\ldots,L$
Drawn $g_{\ell}(\omega_{i})-g_{\ell-1}(\omega_{i})$ for $i=1\ldots m_{\ell}$ and $\ell=0\ldots L$
Estimate convergence rate of $\mathbb{E}\left[g_{\ell}(\omega_{i})-g_{\ell-1}(\omega_{i})\right]$, $\mathbb{V}\left[g_{\ell}(\omega_{i})-g_{\ell-1}(\omega_{i})\right]$ for each $\ell$
Output: Averaged estimated convergence rates $\alpha$ and $\beta$, complexity rate $\gamma$, constants $c_{1},c_{2},c_{3}$
@@
Estimate $L$ as in (40)
Compute $m_{\ell}$, $\ell=0,\ldots,L$ as in (41)
Compute $\mathbb{E}\left[g_{\ell}(\omega_{i})-g_{\ell-1}(\omega_{i})\right]$, for every $\ell$ and $i=1\ldots m_{\ell}$
Build the telescopic sum as in (22).
Output: the estimated mean value.